OpenStudy (anonymous):

Find the surface area when the indicated portion of the curve is rotated about the specified axis. y=cos2x 0<=x<=pi/6 x-axis ill rewrite in first post.

OpenStudy (anonymous):

\[y=\cos2x\] \[0 \le x \le \frac{ \pi }{6 }\]

OpenStudy (anonymous):

about the x-axis

OpenStudy (anonymous):

\[y \prime = -2\sin2x\] \[\large \int\limits_{0}^{\frac{ \pi }{6 }}2 \pi \cos2x \sqrt{1+(-2\sin2x)^{2}} dx\] \[\large \int\limits_{0}^{\frac{ \pi }{6 }}2 \pi \cos2x \sqrt{1+(4\sin^{2}2x)} dx\] and then from here i am stuck

OpenStudy (anonymous):

@amistre64